# VectorFit : Adaptive Singular & Bias Vector Fine-Tuning of Pre-trained Foundation Models

### v1

@@ 5.5 Image Generation
Table 5 summarizes the results of personalized image generation using Dreambooth-style fine-tuning (Ruiz et al., 2023). The evaluation is conducted using three metrics: DINO, CLIP-I, and CLIP-T, as outlined in (Ruiz et al., 2023). Our method recovers an average DINO, CLIP-I, and CLIP-T score of 98.2% of Full-FT accuracy as opposed to 97.3% achieved by LoRA with 0.04% of trainable parameters. Figure 11 provides a visual comparison between Full-FT and VectorFit.
@@ 6.2 Rank Analysis
Proposition 2. Singular vector updates lead to high-rank incremental matrices.
Ignoring the bias, the forward pass of VectorFit can be written as
$y=U\Sigma V^{T}x$ (10)
where $x$ is the input for the current layer and $y$ is the output of the layer before activation. We find the gradient of $\Sigma$ with respect to loss $L$ ($\nabla_{\Sigma}L$) during back-propagation, similar to (Ionescu et al., 2016) and (Hao et al., 2025):
$\nabla_{\Sigma}L=(U^{T}V)\frac{\partial L}{\partial y}x^{T}$ (11)
Here, $\frac{\partial L}{\partial y}$ is the partial derivative of $L$ with respect to $y$, which can also be called as incoming-gradient. $U$ and $V$ are orthogonal matrices and hence $U^{T}V$ is also an orthogonal matrix. We interpret $\nabla_{\Sigma}L$ in Eq. 11 as orthogonal projection of input $x$ transformed by incoming-gradient. Since orthogonal projection is rank-preserving, we hypothesize that the gradients of $\Sigma$ are high-rank (and is a function of the input and incoming-gradient).
Proof. We empirically prove our hypothesis using the singular values of resultant incremental matrix $\Delta^{*}$ for the DeBERTaV3-base model fine-tuned on the COLA dataset using Full-FT and VectorFit. In Full-FT, $\Delta^{*}$,
$\Delta^{*}=W_{init}-W_{final}$ (12)
where $W_{init}\in\mathbb{R}^{d_{r}\times d_{c}}$ and $W_{final}\in\mathbb{R}^{d_{r}\times d_{c}}$ represent the pre-trained and fine-tuned weight matrices, respectively. For VectorFit, $\Delta^{*}$,
$\Delta^{*}=W_{init}-U\Sigma_{final}V^{T}$ (13)
with $U$ and $V$ as the left and right singular matrices of $W_{init}$, and $\Sigma_{final}$ as the singular vector after complete fine-tuning. Singular values of $\Delta^{*}$ from randomly selected layers are plotted in Figure 9 of Appendix D.4. The plots reveal that in Full-FT, a few singular values are significantly larger, while others are relatively small but non-zero, indicating that the incremental matrices are not purely low-rank.
Low-rank adaptation methods operate under the assumption that $\Delta^{*}$ consistently has a rank $r<<min(d_{r},d_{c})$. As a result, their gradient updates are restricted to a smaller subspace compared to Full-FT. The plots show that VectorFit successfully overcomes this drawback and is capable of achieving a full ranked $\Delta^{*}$ with rank $r=min(d_{r},d_{c})$ while also achieving a higher parameter efficiency. In addition, training the bias vectors gives the model an extra degree of freedom to capture the distribution of the fine-tuning dataset.
We can also note that the singular values corresponding to VectorFit is higher than that of Full-FT for several layers. This means that the $\Delta^{*}$ of VectorFit has a higher concentrated energy and variance. This leads to a greater scaling and overall impact on the input vectors which helps in learning stronger internal representations for the data distribution.
@@ 7 Conclusion
Broader Impact
Our approach facilitates the efficient customization of foundational models, which carries potential societal benefits and risks. By reducing the computational and parameter footprint, our method makes customization more cost-effective and accessible.
@@ Appendix A Memory Usage
As discussed in Section 3.1, while retaining the left and right singular matrices increases the overall parameter count, the practical impact is minimal. Figure 5 compares the GPU memory usage of VectorFit and LoRA (r = 1) on the MNLI dataset using DeBERTaV3-base with full precision training. The figure demonstrates that both methods use comparable amounts of memory, where VectorFit requires approximately 200MB of additional memory using 0.08% of trainable parameters. This experiment was conducted on an Nvidia Titan XP GPU with 12GB of RAM.
Figure 5: PyTorch memory trace (Shi & DeVito, ) comparison of 4 training steps for LoRA (r = 1) on the top and VectorFit on the bottom.

### v2

@@ 5.5 Image Generation
Table 6 summarizes the results of personalized image generation using Dreambooth-style fine-tuning [34]. The evaluation is conducted using three metrics: DINO, CLIP-I, and CLIP-T, as outlined in [34]. Our method recovers an average DINO, CLIP-I, and CLIP-T score of 98.2% of Full-FT accuracy as opposed to 97.3% achieved by LoRA with 0.04% of trainable parameters. Figure 12 provides a visual comparison between Full-FT and VectorFit.
@@ 6.2 Rank Analysis
Proposition 2: If the rank of a matrix is high, it contains more information compared to its low-rank counterpart.
Proof. Let $A$ be a matrix of dimensions $m\times n$ with rank $p$. Suppose $A$ is decomposed using Singular Value Decomposition (SVD) as:
$A=U\Sigma V^{T},$
where $U$ and $V$ are orthonormal matrices, and $\Sigma$ is a diagonal matrix containing singular values.
Now, consider a lower-rank approximation of $A$, denoted as $A_{k}$, with rank $k$ such that $k<p$. The truncated SVD of $A_{k}$ is given by:
$A_{k}=U_{k}\Sigma_{k}V_{k}^{T},$
where $U_{k}$, $V_{k}$, and $\Sigma_{k}$ correspond to the top $k$ singular components of $A$.
The difference in Frobenius norm between $A$ and $A_{k}$ is given by:
$\|A-A_{k}\|_{F}=\sqrt{\sum_{i=k+1}^{p}\sigma_{i}^{2}},$
where $\sigma_{i}$ are the singular values of $A$.
This difference represents the information loss due to low-rank approximation. Since the Frobenius norm quantifies the variance within the matrix, reducing the rank leads to a loss in variance information. Therefore, a higher-rank matrix retains more information than its lower-rank approximation.
Proposition 3: Singular vector updates lead to high-rank incremental matrices.
Proof. Consider the incremental weight matrix $\Delta^{*}$:
$\Delta^{*}=W_{\text{init}}-W_{\text{final}}$
Expressing in terms of singular value decomposition:
$\Delta^{*}=U\Sigma_{\text{init}}V^{T}-U\Sigma_{\text{final}}V^{T}$
$\Delta^{*}=U\Delta\Sigma V^{T}$
Since $U$ and $V$ are orthogonal matrices, the rank of $\Delta^{*}$ is upper bounded by the rank of $\Delta\Sigma$. This ensures that $\Delta^{*}$ achieves a full-rank incremental update when all singular values are updated, meaning no singular values are truncated.
Furthermore, by the Eckart–Young–Mirsky theorem [41], the best rank-$k$ approximation of a matrix $A$ is obtained by retaining the top $k$ singular values. Conversely, if no singular values are truncated (i.e., all singular values are updated or remain nonzero), the approximation $\Delta^{*}$ preserves the full rank:
$r=\min(d_{r},d_{c}).$
where $d_{r}$ and $d_{c}$ are the dimensions of weight matrix. This confirms that singular vector updates lead to high-rank incremental matrices. We experimentally verify this by plotting the singular values of $\Delta^{*}$ from randomly selected layers as shown in Figure 10 of Appendix D.4.
Table 7: Ablation study to analyze the efficacy of AVF. The table presents results for L1 regularization, Random Vector Freezing, and AVF.
Method SST-2 CoLA RTE MRPC STSB SQuAD v1.1 SQuAD v2.0
L1 regularization 90.02 64.74 69.38 74.77 84.53 71.3 / 74.5 60.9 / 63.1
Random vector freezing 94.16 70.13 82.82 91.26 90.17 85.1 / 91.7 82.8 / 85.9
Adaptive vector freezing (AVF) 96.10 70.94 84.12 92.16 91.76 87.0 / 93.2 84.4 / 87.6
@@ 6.3 Ablations on Choice of Vectors and AVF
6.4 Efficacy of AVF Against Other Possible Approaches
In this section, we present the experiments conducted by replacing AVF with other possible approaches. We explore L1 regularization and Random vector freezing, where we randomly freeze trainable vectors. From Table 7 it can be seen that AVF consistently outperforms the other two possible approaches. A key distinction between AVF and L1 regularization lies in the granularity at which they operate. Standard L1 regularization regularizes individual singular values, whereas AVF targets the overall strength of training of singular vectors and bias vectors. We hypothesize that regularizing individual singular values may constrain the model’s expressiveness, as it indirectly limits the flexibility of the associated left and right singular vectors. Random vector freezing performs better than L1 regularization but it leads to slower and unstable training. On the other hand, AVF leads to higher performance with stable training.
@@ 7 Conclusion
Appendix
@@ Appendix A Memory Usage
As discussed in Section 3.1, while retaining the left and right singular matrices increases the overall parameter count, the practical impact is minimal. Figure 5 compares the GPU memory usage of VectorFit and LoRA (r = 1) on the MNLI dataset using DeBERTaV3-base with full precision training. The figure demonstrates that both methods use comparable amounts of memory, where VectorFit requires approximately 200MB of additional memory using 0.08% of trainable parameters. Figure 6 shows the GPU memory consumption comparison between LoRA, VectorFit, and SVFT with different base models. We observe that VectorFit consumes higher memory compared to LoRA on larger models while using significantly less memory than that of SVFT. The experiments were conducted on an Nvidia A100 GPU with 40GB of RAM.
Figure 5: PyTorch memory trace [35] comparison of 4 training steps for LoRA (r = 1) on the top and VectorFit on the bottom.
Figure 6: Memory usage.